OpenStudy (anonymous):

- 5(2x+2) - 10x - 17 - 5(2x+2) - 10x - 17 @Mathematics

hero (hero):

Chrissy?

OpenStudy (anonymous):

you know you gotta distribute that -5 into the parenthese first

OpenStudy (anonymous):

-10x-10-10x-17 Now collect like terms: -20x-27 And we are done

OpenStudy (anonymous):

-10x-10-10x-17=???????????????????????????

OpenStudy (anonymous):

-27-20x

OpenStudy (anonymous):

I'm sorry the question is - 5(2x+2) \[\ge\] - 10x - 17

OpenStudy (anonymous):

-20x-27

OpenStudy (anonymous):

-10x-10\[\ge\]-10X-17

OpenStudy (anonymous):

this ques is wrong it will cancel 10x on both sides

OpenStudy (anonymous):

10x - 10 -7x\[\ge\] 3x - 2